find the slope and y intercept of the line in the given equation
y=6/7x-4

Answer by stanbon(75887)

find the slope and y intercept of the line in the given equation
y=6/7x-4
----
y-int = f(0) = (6/7)*0-4 = -4
--------------
slope = (6/7)
==================
Cheers,
Stan H.
==============